OpenStudy (ayonnaleflore):

will fan and medal: Which of the following is not a perfect square trinomial? 169 – 26y + y*2 81 + 18y + y^2 64 + 8y + y^2 25 + 10y + y^2

OpenStudy (welshfella):

you need to find the square root of the first number and check if the coefficient of the term in x = 2* this square root ( note the root may be positive or negative) If this is the case then we have a perfect square.

OpenStudy (welshfella):

OH - and the last term has to be y^2

OpenStudy (ayonnaleflore):

so my answer would be A??

OpenStudy (welshfella):

I notice that its y*2 in the first choice. Could that be a typo?

OpenStudy (ayonnaleflore):

yes its supposed to be y^2

OpenStudy (welshfella):

if its not a typo then one answer is A. But there's another one which id not a perfect square. Try the square root test on B , C and D.

OpenStudy (welshfella):

For B for example square root of 81 is 9 and 2*9 = 18 ( in '18x') so that is a perfect square trinomial Get the idea?

OpenStudy (welshfella):

Try C what is the square root of 64?

OpenStudy (ayonnaleflore):

8

OpenStudy (ayonnaleflore):

8x2=16

OpenStudy (welshfella):

ok so have we got a perfect square here?

OpenStudy (ayonnaleflore):

yes

OpenStudy (welshfella):

check 16 against the number thats in front of the term in x

OpenStudy (ayonnaleflore):

do you mean divide them????? 64/16=4

OpenStudy (welshfella):

No what number multiplies x inr multiplies y in 64 + 8y + y^2 Sorry i've been writing x all the time when i should have written y!!.

OpenStudy (welshfella):

the number in fron of the y

OpenStudy (welshfella):

Its 8 right? (8y)

OpenStudy (welshfella):

sorry i've confused you

OpenStudy (ayonnaleflore):

its fine .......8*16 = 128????

OpenStudy (welshfella):

No C is not a perfect square because there is not a 16 infront of the middle term (the term ending in y) - its 8y . If it was a perfect square it would be 16y.

OpenStudy (ayonnaleflore):

ok so the answer is A???

OpenStudy (welshfella):

No the answer is C We are looking for a triniomial which is not a perfect square. A is a perfect square because SQuare root of 169 = -13 and 2*-13 = -26 which is the number in front of the term in y (middle term)

OpenStudy (ayonnaleflore):

oh......... ok i get it now thank you

OpenStudy (welshfella):

good yw A will expand to ( y - 13)(y - 13) - if we expand that we''ll get y^2 - 26y + y^2

OpenStudy (welshfella):

* A will factor to

OpenStudy (ayonnaleflore):

ok got it